Hobart 26m

Disk VSN: HOB+1005/16000/1024 | Data volume at beginning: 6753.637 GB

  • 1800UT Experiment started okay (Jay).
  • 1820UT Drives broke. Tried to reset drives few times but still no success. Called Warren, he's driving to the telescope site to fix it. (Jay)
  • 2100 UT Tripped limit switch mechanical fault stopped operation. Had to wait for daylight to find (warren)
  • 2230 UT Fixed two limit switches. Found hydraulic leak needs fixing very soon, but schedule can continue. (w)
  • 2243 UT schedule resumed (w)
  • 2300 UT schedule halted something is still bad. suspect main x-bearing in east caused limit switch fault (w)

Yarragadee 12m

Disk VSN: HOB+0106/16000/1024| Data volume at beginning: 4006.655 GB

  • 1800UT Experiment started okay (Jay).

Katherine 12m

Disk VSN: Flexbuffke

  • 1800UT Experiment started late. DBBC3 power levels were bad, had to power-cycle it. still bad even after power cycling. Nothing else I can do atm so putting it on schedule. Recording starts from 1843UT (Jay).
  • Disk-9 is missing in flexbuffke (Jay)
  • 2348 - Schedule halted for network maintenance. (JMc)
  • 0014 - Back on schedule, first good data (JMc)

Hobart 12m

Disk VSN: Flexbuf-2hb

  • 1800UT Experiment started okay (Jay).
  • only 22 out of 36 disks available in flexbuf-2hb for recording at the start of the experiment (Jay).